OpenStudy (anonymous):

\[\frac{ -2n-2 }{ 3 }=12\] @bibby

OpenStudy (solomonzelman):

multiply both sides times 3, add 2 to both sides, divide both sides by -2. OR factor out of -2 on top, multiply by 3/(-2) on both sides,

OpenStudy (anonymous):

i dont get mathematical terms

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle \frac{-2n-2}{3}=12 }\) lets do one step at a time. \(\large\color{slate}{ \displaystyle \frac{-2n-2}{3} \color{blue}{\times 3}=12 \color{blue}{\times 3} }\) multiplied both sides times -3, what do you get?

OpenStudy (anonymous):

i dont really understand when you explain it with words instead of numbers. do you know what im saying

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle \frac{-2n-2}{3} =12 }\) \(\large\color{slate}{ \displaystyle \frac{-2n-2}{3} \color{blue}{\times 3}=12 \color{blue}{\times 3} }\) \(\large\color{slate}{ \displaystyle ?~~~=~~~~? }\)

OpenStudy (solomonzelman):

don't quite get what you mean by saying that you don't get a word explanation. I think I posted pretty nicely the (multiplication times) blue 3's.

OpenStudy (anonymous):

Not that the first one you said.

OpenStudy (solomonzelman):

The first reply you don't get? So I am trying to take it down with you peace by peace. Do you get what I recently said? \(\large\color{slate}{ \displaystyle \frac{-2n-2}{3} \color{blue}{\times 3}=12 \color{blue}{\times 3} }\) (it was this)